OpenStudy (anonymous):

Can somebody help at all with: The Human Memory Models approximate the percentage of information the average person can recall after a certain period of months have passed. I(t) = 75-6ln(t+1) How much does a person retain initially? How much after 24 months? How long does a person retain at least 60% Please help have assignment due

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks like no one wants to do this huh?

OpenStudy (misty1212):

\[f(t)=75-6\ln(t+1)\] How much does a person retain initially? put \(t=0\) and get \[f(0)=75-6\ln(0+1)=75\] since \(\ln(1)=0\)

OpenStudy (misty1212):

oops \[f(24)=75-6\ln(25)\] looks better still need a calculator

OpenStudy (anonymous):

ok have that one, what about the final one would it be 60=75-6ln(t+1)

OpenStudy (misty1212):

solve most of it using algebra

OpenStudy (misty1212):

\[-15=-6\ln(t+1)\] is the first step, then \[\frac{5}{3}=\ln(t+1)\]

OpenStudy (misty1212):

that makes \[t+1=e^{\frac{5}{3}}\]

OpenStudy (misty1212):

subtract one and you are done

OpenStudy (anonymous):

have I put 60% in the correct place of the equation?

OpenStudy (misty1212):

yes the function you are told is the percent, so your equation was correct

OpenStudy (anonymous):

Can I ask you something else please

OpenStudy (anonymous):

Use logarithm laws to re-write R=2/3logbase10(K)-0.9 is the formula for the richter scale and I am supposed to rewrite in the for R=logbase10 (aK^m) where a and m are constants

OpenStudy (misty1212):

can you use the equation tool? or just rewrite it without writing "base ten" ?

OpenStudy (anonymous):

equation tool?

OpenStudy (misty1212):

let me try \[\frac{2}{3}\log(K^{-0.9})\]?

OpenStudy (misty1212):

is that is? or something else?

OpenStudy (anonymous):

LOL I have no idea, where would the a be?

OpenStudy (misty1212):

right below where the little blue box says \(\Sigma\)Equation

OpenStudy (anonymous):

\[R=2/3\log_{10}(K)-0.9 \]

OpenStudy (anonymous):

So I have to rewrite usings log laws to get \[R=\log_{10} (aK ^{m}) \] where a and m are constants derived in re-arrangement

OpenStudy (misty1212):

on the number is outside bring the coefficient inside the log as an exponent

OpenStudy (misty1212):

\[\log(K^{\frac{2}{3}})\]

OpenStudy (misty1212):

ooh you want to bring the \(0.9\) inside as well right?

OpenStudy (anonymous):

\[R=2/3K ^{-0.9}\]

OpenStudy (anonymous):

does that look right?

OpenStudy (misty1212):

no

OpenStudy (anonymous):

oh

OpenStudy (misty1212):

\[\frac{2}{3}\log(K)-0.9\] we want to write this as a single log

OpenStudy (misty1212):

start with \[\frac{2}{3}\log(k)-\log(10^{-0.9})\]

OpenStudy (misty1212):

the use the exponent property of the log and get \[\log(K^{\frac{2}{3}})-\log(10^{-0.9})\]

OpenStudy (misty1212):

and finally use the division property to finish it as \[\log(\frac{K^{\frac{2}{3}}}{10^{-0.9}})\]

OpenStudy (misty1212):

or better still \[\huge \log(10^{0.9}K^{\frac{2}{3}})\]

OpenStudy (anonymous):

can I ask why you minus the log10^-0.9

OpenStudy (misty1212):

because your goal was to bring the number \(-.9\) inside the log

OpenStudy (misty1212):

you have to use \[\log(A)-\log(B)=\log(\frac{A}{B})\]so you need to write the number \(-0.9\) as a logarithm

OpenStudy (anonymous):

\[\log_{10} ^{-0.9}\]

OpenStudy (misty1212):

yeah that one

OpenStudy (anonymous):

or \[\log10^{-0.9}\]

OpenStudy (misty1212):

\[-0.9=\log(10^{-0.9})\]

OpenStudy (anonymous):

You are a treasure, thank you so very very much xoxo

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]